OpenStudy (anonymous):

find f'(x) of f(x)= 4x + 5sin^4x

OpenStudy (anonymous):

f'(x) = 4 + 20cos4x chain rule used to differentiate the second term

OpenStudy (anonymous):

i dont see how that is the answer

OpenStudy (anonymous):

well derivative of 4x = 4 right?

OpenStudy (anonymous):

something would be to the third power though, wouldnt it?

OpenStudy (anonymous):

that x is not in the power \[f(x) = 4x +5\sin^4 x\] i dont know if that makes a difference

OpenStudy (anonymous):

standard result: if f(x) = ax^n then f'(x) = an x^(n-1) so for 4x = 4x^1 f'(x) = 4*1 x^(1-1) = 4 x^0 = 4

OpenStudy (anonymous):

oh sorry i misread the second term - i took it to be 5sin4x

OpenStudy (anonymous):

so would it be 4+20sinx^3 * cosx?

OpenStudy (anonymous):

ok derivative of second term is 4 sin^3 x * 5 cos x = 20cosxsin^3 x 4 + 20cosxsin^3x is your answer

OpenStudy (anonymous):

yes you are right